OpenStudy (anonymous):

solve the differential equation y'=3y(x+1)^2/(y-1)

OpenStudy (anonymous):

Assuming y'=dy/dx, you can use separation of variables to get [(y-1)dy]/(3y) = (x+1)^2dx. Then just integrate both sides. Don't forget the constant of integration.